# MetaGS: A Meta-Learned Gaussian-Phong Model for Out-of-Distribution 3D Scene Relighting

### v2

@@ C.2 Clarification on Meta-learning
      Meta-learn       w/o Meta-Learn       Larger batch (×$\times m$)
@@ C.3 Multi-light Relighting
To validate MetaGS’s capability in handling complex lighting conditions, we conduct experiments with test sets containing 22223333 light sources. As shown in Figure 8, our method demonstrates robust generalization beyond single-light scenarios.
@@ C.5 In-distribution Novel View Relighting
This experiment follows the setup in NRHints [33], where both training and testing lights are randomly located in the upper hemisphere surrounding the scene). Each view is illuminated by a single point light placed at a unique position, though all light positions are distributed within the same region. During the testing phase, we evaluate on novel viewpoints under new, in-distribution light positions. Similar to the OOD setting, we generate $600$ OLAT images per synthetic scene using Blender ($500$ for training and $100$ for testing). For real-world scenes, we use only $600$ training images–—a significantly smaller subset compared to the full NRHints dataset.
@@
3DGS [9] 19.55 0.9130 0.0580 21.13 0.9417 0.0470 23.22 0.9248 0.0796 21.30 0.9265 0.0615
Relightable 3DGS [7] 19.62 0.9088 0.0610 21.07 0.9390 0.0502 23.20 0.9224 0.0838 21.30 0.9234 0.0650
GaussianShader [8] 19.62 0.9137 0.0586 20.92 0.9394 0.0500 23.12 0.9241 0.0833 21.22 0.9257 0.0640
NRHints [33] 18.99 0.9087 0.0731 35.81 0.9847 0.0281 36.64 0.9637 0.0481 30.48 0.9524 0.0498
WildLight [5] 22.71 0.9284 0.0499 23.28 0.9068 0.1018 24.65 0.9271 0.0737 23.55 0.9208 0.0751
GS3 [1] 28.20 0.9659 0.0396 29.53 0.9755 0.0317 29.34 0.9520 0.0618 29.02 0.9645 0.0444
@@ Appendix D Preliminaries
Our method builds upon 3D Gaussian Splatting (3DGS) [9], which employs explicit Gaussian points for 3D scene representation. Each Gaussian’s geometry is characterized by an opacity $o\in[0,1]$, a center point $\mu\in\mathbb{R}^{3\times 1}$ and a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$:
@@ D.2 Blinn-Phong Reflection Model
The Phong model is widely adopted in computer graphics for simulating the interaction of light with object surfaces [18]. In this model, the reflected light transport on a surface has three components: ambient ($L_{a}$), diffuse ($L_{d}$), and specular reflections ($L_{s}$). The ambient reflection component represents the constant illumination present in the environment, simulating how light scatters and reflects off other surfaces to establish a baseline brightness level. The diffuse reflection component, governed by Lambertian diffusion law, illustrates the scattering of light in multiple directions upon striking a rough surface, leading to a matte appearance. This component is calculated based on the angle between the light direction ($\mathbf{l}$) and the surface normal ($\mathbf{n}$), ensuring that surfaces facing the light source appear brighter. The specular term is computed based on the angle between the light direction vector and the bisector ($\mathbf{h}$) between the view direction ($\mathbf{v}$) and the light direction ($\mathbf{l}$). The Phong model also incorporates a shininess exponent that governs the spread of the specular highlight, allowing the model to represent the surface’s glossiness. We select the Blinn-Phong model [2] to provide the illumination decomposition priors to 3DGS for its simple yet strong physical prior and computation efficiency. With little addition to the Gaussian’s attribute, our model produces realistic lighting effects.
@@ Appendix E Hyperparameters
      learning rate       $\beta$       same as 3DGS [9]
@@
      meta-train task number       $m$       $5$

### v3

@@ C.2 Clarification on Meta-learning
  Meta-learn   w/o Meta-Learn   Larger batch ($\times m$)
@@ C.3 Multi-light Relighting
To validate MetaGS’s capability in handling complex lighting conditions, we conduct experiments with test sets containing 2233 light sources. As shown in Figure 8, our method demonstrates robust generalization beyond single-light scenarios.
@@ C.5 In-distribution Novel View Relighting
This experiment follows the setup in NRHints Zeng et al. (2023), where both training and testing lights are randomly located in the upper hemisphere surrounding the scene). Each view is illuminated by a single point light placed at a unique position, though all light positions are distributed within the same region. During the testing phase, we evaluate on novel viewpoints under new, in-distribution light positions. Similar to the OOD setting, we generate $600$ OLAT images per synthetic scene using Blender ($500$ for training and $100$ for testing). For real-world scenes, we use only $600$ training images–—a significantly smaller subset compared to the full NRHints dataset.
@@
3DGS Kerbl et al. (2023) 19.55 0.9130 0.0580 21.13 0.9417 0.0470 23.22 0.9248 0.0796 21.30 0.9265 0.0615
Relightable 3DGS Gao et al. (2023) 19.62 0.9088 0.0610 21.07 0.9390 0.0502 23.20 0.9224 0.0838 21.30 0.9234 0.0650
GaussianShader Jiang et al. (2024) 19.62 0.9137 0.0586 20.92 0.9394 0.0500 23.12 0.9241 0.0833 21.22 0.9257 0.0640
NRHints Zeng et al. (2023) 18.99 0.9087 0.0731 35.81 0.9847 0.0281 36.64 0.9637 0.0481 30.48 0.9524 0.0498
WildLight Cheng et al. (2023) 22.71 0.9284 0.0499 23.28 0.9068 0.1018 24.65 0.9271 0.0737 23.55 0.9208 0.0751
GS3 Bi et al. (2024) 28.20 0.9659 0.0396 29.53 0.9755 0.0317 29.34 0.9520 0.0618 29.02 0.9645 0.0444
@@ Appendix D Preliminaries
Our method builds upon 3D Gaussian Splatting (3DGS) Kerbl et al. (2023), which employs explicit Gaussian points for 3D scene representation. Each Gaussian’s geometry is characterized by an opacity $o\in[0,1]$, a center point $\mu\in\mathbb{R}^{3\times 1}$ and a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$:
@@ D.2 Blinn-Phong Reflection Model
The Phong model is widely adopted in computer graphics for simulating the interaction of light with object surfaces Phong (1998). In this model, the reflected light transport on a surface has three components: ambient ($L_{a}$), diffuse ($L_{d}$), and specular reflections ($L_{s}$). The ambient reflection component represents the constant illumination present in the environment, simulating how light scatters and reflects off other surfaces to establish a baseline brightness level. The diffuse reflection component, governed by Lambertian diffusion law, illustrates the scattering of light in multiple directions upon striking a rough surface, leading to a matte appearance. This component is calculated based on the angle between the light direction ($\mathbf{l}$) and the surface normal ($\mathbf{n}$), ensuring that surfaces facing the light source appear brighter. The specular term is computed based on the angle between the light direction vector and the bisector ($\mathbf{h}$) between the view direction ($\mathbf{v}$) and the light direction ($\mathbf{l}$). The Phong model also incorporates a shininess exponent that governs the spread of the specular highlight, allowing the model to represent the surface’s glossiness. We select the Blinn-Phong model Blinn (1977) to provide the illumination decomposition priors to 3DGS for its simple yet strong physical prior and computation efficiency. With little addition to the Gaussian’s attribute, our model produces realistic lighting effects.
@@ Appendix E Hyperparameters
  learning rate   $\beta$   same as 3DGS Kerbl et al. (2023)
@@
  meta-train task number   $m$   55